⦁   Suppose you are the manager of a restaurant and contemplating on purchasing a new machine that will cost 200,000 AED and has a useful life of five years. The machine will yield (year-end) cost reductions to of 30,000 in year 1, 40,000 in year 2; 50,000 in year 3; 60,000 in year 4 and 70,000 in year 5

⦁   What is the present value of the cost saving of the machine if the interest rate is 7%?


⦁   Should the manager purchase the machine?


⦁   Suppose demand and supply of good X is given by the following equations:

      

⦁   Find the inverse demand function.

⦁   How much consumer surplus do consumers receive when Px = 30 AED?

⦁   Find the inverse supply function.


⦁   How much surplus do producers receive when Px = 50 AED?


⦁   What are the equilibrium quantity and price in this market?


⦁   Determine the quantity demanded, the quantity supplied, and the magnitude of surplus if a price floor of 45 AED is imposed in this market.

⦁   Determine the quantity demanded, the quantity supplied, and the magnitude of shortage if a price ceiling of 32 AED is imposed in this market. Also, determine the full economic price paid by the consumers and the nonpecuniary price.
Answer #1

Future cost reductions are future benefits need to be discounted

Present worth of Benefits=30000/(1.07)+40000/(1.07^2)+50000/(1.07^3)+60000/(1.07^4)+70000/(1.07^5)

Present worth of cost=200000

Present worth of investment = Present worth of bemefits-Present worth of Costs=199472.57-200000=-527.43

As present worth of project is negative manager should not option for project.